# The Security Threat of Compressed Projectors in Large Vision-Language Models

### v1

@@ 4.1 Experimental Settings
Datasets. We randomly sampled 2000 image-question pairs from VQA v2 Antol et al. (2015) to construct our primary dataset, referred to as “VQA v2 2000”. Additionally, we include results for other datasets in Sec. 4.7, including ImageNet Deng et al. (2009), VizWiz Bigham et al. (2010), and COCO Lin et al. (2014).
@@
Adversarial attack. We generated adversarial perturbation on clean images using PGD-$l_{\infty}$ Madry et al. (2018) and CW-$l_{2}$ Carlini and Wagner (2017) to generate adversarial noise on clean images. For PGD-$l_{\infty}$, we set the attack step to 2/255, maximum perturbation to 8/255, and the number of attack steps to 20. For CW-$l_{2}$, we configured the attack step to 0.1 (equivalent to 2.55/255), with a constant of 0.005 and zero confidence.
Training details. (1) Surrogate compressed VLPs: we followed BLIP-2’s two-stage pre-training process Li et al. (2023). In the first stage, vision-language representation learning was performed using a frozen VE. In the second stage, both the VE and LLM were kept frozen while performing vision-to-language generative learning. To ensure architectural diversity, we used different LLMs during pre-training (e.g., opt-2.7B and opt-6.7B) compared to the target LVLM. Pre-training was conducted using the COCO datasets.
(2) Surrogate uncompressed VLPs: we followed LLAVA’s official pre-training and fine-tuning settings. Similar to our compressed models, we used a different LLM during pre-training and fine-tuning to align with our gray-box experimental setup.
@@ 4.2 Results of White-box Setting
The results of adversarial attacks under the white-box setting are summarized in Tab. 1. Here, $\mathcal{L}_{\text{VE}}$ represents an adversarial attack specifically targeting the VE, as formally defined in Eq. 2, while $\mathcal{L}_{\text{VLP}}$ denotes an attack directed at the VLP, defined in Eq. 3. A key observation emerges under this setting: for compressed projectors, adversarial attacks targeting VLP achieve superior performance compared to those targeting VE. This discrepancy underscores the heightened security vulnerability of compressed projection. Conversely, in the case of uncompressed projectors, attacks against VLP yield results comparable to those targeting VE, demonstrating the enhanced robustness of uncompressed projectors. These initial findings from white-box attack experiments reveal critical security implications for compressed projector.
@@
[…]
@@ 4.3 Results of Attacks via Surrogate VLPs from Other LVLMs
We employ the surrogate VLP from other similar LVLMs to launch attacks against our target models. Specifically, for compressed projectors, we conduct experiments using InstructBLIP models that have been trained with various LLMs. For uncompressed projectors, we utilize LLaVA v1.5 as a surrogate model to attack LLaVA v1.6. As demonstrated in Tab. 2, our results indicate that uncompressed projectors maintain superior robustness under gray-box attack conditions, whereas attacks targeting compressed projectors lead to further degradation in model performance.
@@ 4.4 Results of Attacks via Surrogate VLPs Trained from Scratch
We further explored training the surrogate VLP from scratch, using a different LLM than the target LVLM within the context of our gray-box attack scenario. Our experimental results, as evidenced by Tab. 3, demonstrate consistent findings with those presented in Tab. 2: uncompressed projectors exhibit superior security performance even under more challenging attacking scenarios.
@@
[…]
@@ 4.5 Is Security Dependent on Architecture or the Quantity of Visual Tokens?
Our experimental results, as summarized in Tabs. 1, 2 and 3, demonstrate that compressed projectors exhibit significant security vulnerabilities compared to their uncompressed counterparts. However, it is important to note that uncompressed projectors generally possess a larger number of tokens (e.g., the MLP employed in our laboratory utilizes 576 visual tokens, whereas the Q-former operates with only 32 visual tokens). This raises an essential question: does the reduced security of the Q-former stem from the inherent characteristics within its architecture, or merely from its smaller quantity of visual tokens?
@@
We then conduct adversarial attacks on LLaVA-v1.5-7B with varying numbers of visual tokens. The results in Tab. 4 reveal that even when only 36 visual tokens are utilized (comparable to Q-former’s), the attack performance on VLP does not surpass that on VE, implying its robustness against attacks. This finding conclusively demonstrates that the enhanced robustness of uncompressed projectors is not merely dependent on the quantity of visual tokens, but rather arises inherently from their structural design. Furthermore, this implies that in pursuit of efficiency, one can opt for an uncompressed projector while employing techniques such as spatial pooling to minimize its visual markers. Even when their respective quantities of visual tokens are comparable in number, the security provided by an uncompressed projector remains markedly superior to that of its compressed ones.
@@ 4.6 Further Strategies to Deteriorate the Security of Compressed Projectors
Table 3 presents the attack via training surrogate VLPs. We employ a single surrogate VLP for simplicity. However, multiple training runs can yield additional surrogate models, which we then leverage collectively to launch attacks against the target model, as illustrated in Fig. 1. The results summarized in Tab. 5 demonstrate that this multi-VLPs attack strategy further degrades the performance of the compressed projector.
@@
Attack $N$ The target models (InstructBLIP) Avg.
[…]
@@
We start by setting $\beta=0$ to specifically examine attacks designed to target VLP, enabling us to analyze whether incorporating VLP improves attack effectiveness. We then combine the $\mathcal{L}_{\text{VE}}$ and $\mathcal{L}_{\text{VLP}}$ losses as Eq. 4. As shown in Tab. 6, $\mathcal{L}_{\text{TCP}}$ achieves superior attack performance compared to other approaches. Additionally, some adversarial images are provided in Appendix A.
Attack $N$ $\beta$ The target models (InstructBLIP) Avg.
[…]
@@ 4.7 Results on Additional Datasets
Using InstructBLIP-based Vicuna-7B, we performed a thorough security evaluation to analyze potential vulnerabilities of compressed projectors across diverse datasets extending beyond VQA v2 in Tab. 7. Our assessment revealed that in a gray-box attack scenario, where only the target model’s VE weights and VLP architecture are known, we were able to significantly degrade model performance, highlighting the inherent security risks associated with compressed projectors.
@@ 4.8 Ablation Study on Pre-Training Tasks
BLIP-2’s first-stage pre-training consists of three tasks: Image-Text Matching (ITM), Image-Text Contrastive Learning (ITC), and Image-Grounded Text Generation (Image Captioning, IC). By default, we adopted the standard BLIP-2 training configuration, employing all three tasks concurrently. However, our experimental results presented in Tab. 8 reveal an important insight: training the surrogate VLP exclusively with the IC task achieves attack performance comparable to utilizing all three tasks together. This not only reduces the overhead associated with training the surrogate model but also amplifies the security risks posed by the compressed projector.
@@
[…]
@@ 5 Conclusion
This study investigates the security vulnerabilities of VLP structures within two representative LVLM frameworks. Our analysis exposes severe security susceptibilities in compressed projectors, while highlighting the robust performance of uncompressed alternatives. Through rigorous empirical evaluation, we demonstrate that this vulnerability originates inherently from the architectural design of the compression encoder itself and remains independent of the visual token quantity. These findings sound a cautionary note regarding the security implications for LVLMs employing compressed projectors, while encouraging researchers to adopt a more comprehensive understanding of VLP performance characteristics.
Limitations
We summarize the limitations of our study as follows: (1) The attack methodology has certain constraints. We implemented the attacks using only basic PGD and CW methods on VQA and image captioning tasks. Notably, we did not employ advanced techniques designed to enhance adversarial transferability, such as DI Xie et al. (2019) and MI Dong et al. (2018). Additionally, we limited our attacks to these two tasks without extending to other domains. Despite this, we successfully exposed the security vulnerabilities of compressed projectors, and incorporating DI and MI methods could potentially amplify the robustness risks associated with these models. (2) Our study focuses exclusively on analyzing the security vulnerabilities of both compressed and uncompressed projectors. However, we do not investigate potential defense mechanisms against the attacks. A discussion of possible defensive strategies would provide valuable insights into enhancing the robustness of compressed projectors against such attacks.
@@ Appendix A Visualized Examples of Attacks
We present some visualized examples from VQA v2 2000 and InstructBLIP(Vicuna-7B) in Fig. 2 and Fig. 3.
@@
  • For each row, the left image is the clean image, the center image is the adversarial image, and the right image is the adversarial perturbation magnified by a factor of 16 (i.e., when perturbation limitation $\epsilon=8$, $-\epsilon=-8\to 0$ and $\epsilon=8\to 255$). The adversarial image is obtained from a clean image plus an adversarial perturbation.
These examples demonstrate that the attack successfully misdirects the LVLM even under the gray-box setting, highlighting security vulnerabilities in compression encoders.
@@
Figure 2: Some adversarial examples of attacks on InstructBLIP(Vicuna-7B).
@@
Figure 3: Some adversarial examples of attacks on InstructBLIP(Vicuna-7B).

### v2

@@ 4.1 Experimental Settings
Datasets. We randomly sampled 2000 image-question pairs from VQA v2 Antol et al. (2015) to construct our primary dataset, referred to as “VQA v2 2000”. Additionally, we include results for other datasets in Sec.˜4.7, including ImageNet Deng et al. (2009), VizWiz Bigham et al. (2010), and COCO Lin et al. (2014).
@@
Adversarial attack. We generated adversarial perturbations on clean images using PGD-$l_{\infty}$ Madry et al. (2018) and CW-$l_{2}$ Carlini and Wagner (2017). For PGD-$l_{\infty}$, we set the attack step to 2/255, maximum perturbation to 8/255, and the number of attack steps to 20. For CW-$l_{2}$, we configured the attack step to 0.01 (equivalent to 2.55/255), with a constant of 0.005 and zero confidence.
Training details. (1) Surrogate compressed VLPs: we followed BLIP-2’s two-stage pre-training process Li et al. (2023). In the first stage, vision-language representation learning was performed using a frozen VE. In the second stage, both the VE and LLM were kept frozen while performing vision-to-language generative learning. To ensure architectural diversity, we used different LLMs during pre-training (e.g., opt-2.7B and opt-6.7B) compared to the target LVLM. Pre-training was conducted using the COCO datasets. (2) Surrogate uncompressed VLPs: we followed LLAVA’s official pre-training and fine-tuning settings. Similar to our compressed models, we used a different LLM during pre-training and fine-tuning to align with our gray-box experimental setup.
@@ 4.2 Results of White-box Setting
The results of adversarial attacks under the white-box setting are summarized in Tab.˜1. Here, $\mathcal{L}_{\text{VE}}$ represents an adversarial attack specifically targeting the VE, as formally defined in Eq.˜2, while $\mathcal{L}_{\text{VLP}}$ denotes an attack directed at the VLP, defined in Eq.˜3. A key observation emerges under this setting: for compressed projectors, adversarial attacks targeting VLP achieve superior performance compared to those targeting VE. This discrepancy underscores the heightened security vulnerability of compressed projection. Conversely, in the case of uncompressed projectors, attacks against VLP yield results comparable to those targeting VE, demonstrating the enhanced robustness of uncompressed projectors. These initial findings from white-box attack experiments reveal critical security implications for compressed projectors.
@@
method surrogate Q-former $\text{FlanT5}_{\text{XL}}$ Vicuna-7B $\text{FlanT5}_{\text{XXL}}$ Vicuna-13B
@@ 4.3 Results of Attacks via Surrogate VLPs from Other LVLMs
We employ the surrogate VLP from other similar LVLMs to launch attacks against our target models. Specifically, for compressed projectors, we conduct experiments using InstructBLIP models that have been trained with various LLMs. For uncompressed projectors, we utilize LLaVA v1.5 as a surrogate model to attack LLaVA v1.6. As demonstrated in Tab.˜2, our results indicate that uncompressed projectors maintain superior robustness under gray-box attack conditions, whereas attacks targeting compressed projectors lead to further degradation in model performance.
@@ 4.4 Results of Attacks via Surrogate VLPs Trained from Scratch
We further explored training the surrogate VLP from scratch, using a different LLM than the target LVLM within the context of our gray-box attack scenario. Our experimental results, as evidenced by Tab.˜3, demonstrate consistent findings with those presented in Tab.˜2: uncompressed projectors exhibit superior security performance even under more challenging attacking scenarios.
@@
method Q-former $\text{FlanT5}_{\text{XL}}$ Vicuna-7B $\text{FlanT5}_{\text{XXL}}$ Vicuna-13B
@@ 4.5 Is Security Dependent on Architecture or the Quantity of Visual Tokens?
Our experimental results, as summarized in Tabs.˜1, 2 and 3, demonstrate that compressed projectors exhibit significant security vulnerabilities compared to their uncompressed counterparts. However, it is important to note that uncompressed projectors generally possess a larger number of tokens (e.g., the MLP employed in our laboratory utilizes 576 visual tokens, whereas the Q-former operates with only 32 visual tokens). This raises an essential question: does the reduced security of the Q-former stem from the inherent characteristics within its architecture, or merely from its smaller quantity of visual tokens?
@@
We then conduct adversarial attacks on LLaVA-v1.5-7B with varying numbers of visual tokens. The results in Tab.˜4 reveal that even when only 36 visual tokens are utilized (comparable to Q-former’s), the attack performance on VLP does not surpass that on VE, implying its robustness against attacks. This finding conclusively demonstrates that the enhanced robustness of uncompressed projectors is not merely dependent on the quantity of visual tokens, but rather arises inherently from their structure design. Furthermore, this implies that in pursuit of efficiency, one can opt for an uncompressed projector while employing techniques such as spatial pooling to minimize its visual markers. Even when their respective quantities of visual tokens are comparable in number, the security provided by an uncompressed projector remains markedly superior to that of its compressed ones.
@@ 4.6 Further Strategies to Deteriorate the Security of Compressed Projectors
Table˜3 presents the attack via training surrogate VLPs. We employ a single surrogate VLP for simplicity. However, multiple training runs can yield additional surrogate models, which we then leverage collectively to launch attacks against the target model, as illustrated in Fig.˜1. The results summarized in Tab.˜5 demonstrate that this multi-VLPs attack strategy further degrades the performance of the compressed projector.
@@
Attack $K$ The target models (InstructBLIP) Avg.
method $\text{FlanT5}_{\text{XL}}$ Vicuna-7B $\text{FlanT5}_{\text{XXL}}$ Vicuna-13B
@@
We start by setting $\beta=0$ to specifically examine attacks designed to target VLP, enabling us to analyze whether incorporating VLP improves attack effectiveness. We then combine the $\mathcal{L}_{\text{VE}}$ and $\mathcal{L}_{\text{VLP}}$ losses as Eq.˜4. As shown in Tab.˜6, $\mathcal{L}_{\text{TCP}}$ achieves superior attack performance compared to other approaches. Additionally, some adversarial images are provided in Appendix˜A.
Attack $K$ $\beta$ The target models (InstructBLIP) Avg.
method $\text{FlanT5}_{\text{XL}}$ Vicuna-7B $\text{FlanT5}_{\text{XXL}}$ Vicuna-13B
@@ 4.7 Results on Additional Datasets
Using InstructBLIP-based Vicuna-7B, we performed a thorough security evaluation to analyze potential vulnerabilities of compressed projectors across diverse datasets extending beyond VQA v2 in Tab.˜7. Our assessment revealed that in a gray-box attack scenario, where only the target model’s VE weights and VLP architecture are known, we were able to significantly degrade model performance, highlighting the inherent security risks associated with compressed projectors.
@@ 4.8 Ablation Study on Pre-Training Tasks
BLIP-2’s first-stage pre-training consists of three tasks: Image-Text Matching (ITM), Image-Text Contrastive Learning (ITC), and Image-Grounded Text Generation (Image Captioning, IC). By default, we adopted the standard BLIP-2 training configuration, employing all three tasks concurrently. However, our experimental results presented in Tab.˜8 reveal an important insight: training the surrogate VLP exclusively with the IC task achieves attack performance comparable to utilizing all three tasks together. This not only reduces the overhead associated with training the surrogate model but also amplifies the security risks posed by the compressed projector.
@@
method $\text{FlanT5}_{\text{XL}}$ Vicuna-7B $\text{FlanT5}_{\text{XXL}}$ Vicuna-13B
@@
4.9 Results on Various Attack Methods
In addition to PGD and CW in Tab.˜3, we also implemented an additional attack method: MI-FGSM Dong et al. (2018). Under gray-box settings, we conducted MI-FGSM attacks and extended the analysis presented in Tab.˜3 of the original paper to incorporate methods beyond PGD and CW. The specific parameter configuration for MI-FGSM included setting the momentum value to 0.9. Our findings remain consistent with the conclusions in Secs.˜4.3 and 4.4. Specifically, uncompressed projectors demonstrate superior robustness under gray-box attack settings, while attacks targeting compressed projectors result in further degradation of model performance.
method MLP Vicuna-7B Mistral-7b Vicuna-13B Hermes-Yi-34B
Clean 78.32 79.10 75.61 79.07
MI-FGSM $\mathcal{L}_{\text{VE}}$ 71.65 68.66 72.85 72.01
Vicuna-v1.3-7B 72.40 69.24 73.48 73.10
Vicuna-v1.3-13B 72.90 69.99 73.58 73.49
method Q-former $\text{FlanT5}_{\text{XL}}$ Vicuna-7B $\text{FlanT5}_{\text{XXL}}$ Vicuna-13B
Clean 72.43 75.53 71.86 68.22
MI-FGSM $\mathcal{L}_{\text{VE}}$ 49.99 52.41 49.40 47.06
opt-6.7B 41.71 44.95 41.95 41.54
Table 9: Additional results of Tab.˜3 on MI-FGSM.
@@ 5 Conclusion
This study investigates the security vulnerabilities of VLP structures within two representative LVLM frameworks. Our analysis exposes severe security susceptibilities in compressed projectors, while highlighting the robust performance of uncompressed alternatives. Through rigorous empirical evaluation, we demonstrate that this vulnerability originates inherently from the architectural design of the compressed projector itself and remains independent of the visual token quantity. These findings sound a cautionary note regarding the security implications for LVLMs employing compressed projectors, while encouraging researchers to adopt a more comprehensive understanding of VLP performance characteristics.
6 Discussion of Defense Mechanism
We briefly discuss two potential strategies for enhancing model robustness:
(1) Optimizing uncompressed projector designs for improved efficiency. While uncompressed projectors have demonstrated strong security properties, their inefficiency stems from a large number of visual tokens. However, our experiments in Tab.˜4 reveal that not all visual tokens are essential for performance. By applying 2x2 pooling to LLaVA vision tokens, we successfully reduced the number of visual tokens to 25% of the original amount, resulting in only a marginal 1.3% drop in accuracy, and the LVLMs are still robust. This suggests significant potential for optimizing uncompressed projectors by eliminating redundant visual tokens while retaining their inherent security benefits, thereby enhancing the robustness.
(2) Hybrid architectures combining compressed and uncompressed projectors. Compressed projectors offer high efficiency and sufficient accuracy but lack robustness, whereas uncompressed projectors provide strong robustness and accuracy at the cost of efficiency. Building on insights from Tab.˜4, pooling significantly reduced visual tokens, and the accuracy does not significantly decrease while maintaining safety, we propose a hybrid approach. For instance, we could pool the visual tokens from an uncompressed projector (originally 576 tokens) to 25% of their original count (144 tokens), and supplement this with 32 tokens from a compressed projector, resulting in a total of 144+32=176 tokens. This combination aims to achieve a balance of high accuracy, efficiency, and robustness. Unfortunately, due to challenges in training LVLMs with compressed projectors (lack of fully open-sourced training code of Q-former based LVLMs), we may face limitations in fully implementing this.
7 Limitations
We summarize the limitations of our study as follows: (1) The attack methodology has certain constraints. We implemented the attacks using only basic PGD, CW and MI-FGSM methods on VQA and image captioning tasks. Notably, we did not employ advanced techniques designed to enhance adversarial transferability, e.g., DI-FGSM Xie et al. (2019). Additionally, we limited our attacks to these two tasks without extending to other domains. Despite this, we successfully exposed the security vulnerabilities of compressed projectors, and incorporating DI and MI methods could potentially amplify the robustness risks associated with these models. (2) Our study focuses exclusively on analyzing the security vulnerabilities of both compressed and uncompressed projectors. However, we do not investigate potential defense mechanisms against the attacks. A discussion of possible defensive strategies would provide valuable insights into enhancing the robustness of compressed projectors against such attacks.
8 Acknowledgments
This work was supported by Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), the National Key R&D Program of China (2023YFB4502200), the National Natural Science Foundation of China (No. 62376024, 62325405, 62104128, 62203257, 62031017, 62406159, U21B2031), Tsinghua University Initiative Scientific Research Program, Beijing National Research Center for Information Science, Technology (No. BNR2024TD03001) and Beijing Innovation Center for Future Chips.
@@ Appendix A Visualized Examples of Attacks
We present some visualized examples from VQA v2 2000 and InstructBLIP-Vicuna-7B in Fig.˜2 and Fig.˜3.
@@
  • For each row, the left image is the clean image, the center image is the adversarial image, and the right image is the adversarial perturbation magnified by a factor of 16 (i.e., when perturbation limitation $\epsilon=8$, $\epsilon=-8\to 0$ and $\epsilon=8\to 255$). The adversarial image is obtained from a clean image plus an adversarial perturbation.
These examples demonstrate that the attack successfully misdirects the LVLM even under the gray-box setting, highlighting security vulnerabilities in compressed projectors.
@@
Figure 2: Some adversarial examples of attacks on InstructBLIP-Vicuna-7B.
@@
Figure 3: Some adversarial examples of attacks on InstructBLIP-Vicuna-7B.